# Health Text Simplification: An Annotated Corpus for Digestive Cancer Education and Novel Strategies for Reinforcement Learning

### v2

@@ 4.1 Text Simplification
Table 2 presents the performance on the withheld SimpleDC test set, including three LLMs: Llama-2-7b-chat-hf, Llama-2-7b-32K-Instruct, and GPT4 Turbo. In the zero-shot setting (Exp. 01, 03, and 13), Llama-2-7b-chat-hf achieved the highest SARI performance, although GPT-4 achieved the lowest FKGL. For both Llama 2 models (Exp. 02 and 04), SFT drastically increased the performance of SARI, BLEU, BERTScore, and ROUGE and reduced FKGL by approximately two grade levels. Even with the relatively small SimpleDC training set, SFT allows the Llama 2 models to learn an understanding of the SimpleDC annotation approach and voice of the annotators. SFT performance for the two Llama 2 variants is similar across all metrics, including SARI (47.92 vs. 47.82). The success of the RL/RLHF varies by Llama 2 variant and reward strategy. In the RL/RLHF-only experimentation, Llama-2-7b-32K-Instruct (Exp. 09-11) consistently outperforms Llama-2-7b-chat-hf (Exp. 05-07) across metrics, including higher SARI and lower FKGL. For Llama-2-7b-32K-Instruct, the proposed RLHF, $R_{FKGL+OvS}$ (Exp. 11), increases performance from the zero-shot approach from 27.21 to 59.02 SARI. Combining SFT and RLHF for Llama-2-7b-32K-Instruct (Exp. 12) yields the best-performing system in terms of SARI (74.18), BLEU (0.77) and BERTScore (0.97). For this configuration, the FKGL metric suggests that this approach yields text at a relatively low reading level of 7.95. These results demonstrate that RL/RLHF improve text simplification performance both with and without prior SFT on labeled data; however, combining SFT and RL/RLHF yields the highest performance. Additionally, these results demonstrate the importance of incorporating human feedback through RLHF over RL without human feedback (Exp. 11 vs 10).
Conversely, the GPT4 Turbo model (Exp. 13-17) exhibited lower scores across all adequacy metrics (SARI, BLEU, BERTScore, and ROUGE) compared to the Llama 2 models (Exp. 01-12) but achieved the lowest FKGL. Lower performance does not necessarily indicate lower quality simplifications generated by GPT-4. Based on our qualitative review, GPT-4 edited the original text more extensively, resulting in a larger divergence between the original and simplified text. Additionally, the voice or writing style of GPT-4 appears to differ from that of our annotators, contributing to the lower performance for adequacy metrics. The FKGL-Enhanced prompting strategy (Exp. 15) achieved the best performance across adequacy metrics, and the CoT prompting strategy (Exp. 16) achieved the lowest FKGL. The Self-Correction approach (Exp. 17) achieved a balance between adequacy and reading level at 22.86 SARI and 4.48 FKGL.
@@ 4.2 Domain Adaptation
We explored the potential for RL/RLHF to adapt an out-of-domain text simplification model to a target domain, specifically the adaptation of a model trained on Med-EASi to our SimpleDC corpus. Table 2 (Exp. 18-23) presents the domain adaptation results, where all models were first trained through SFT on the Med-EASi training set [37]. The models were subsequently adapted through RL/RLHF using unlabeled texts related to digestive cancer that are not contained in SimpleDC. This experimentation was intended to evaluate the potential to use external out-of-domain corpora to improve in-domain performance. The performance gains associated with using RL/RLHF for domain adaptation vary by Llama 2 variant and specific reward function. The Llama-2-7B-32K-Instruct with $R_{FKGL+OvS}$ reward (Exp. 23) led to an increase across all adequacy metrics in the in-domain experimentation (comparing with Exp. 21), including increased SARI from 26.70 to 45.71, BLEU from 0.36 to 0.63, and ROUGE 0.52 to 0.78. Simultaneously, the FKGL decreased from 8.25 to 7.50, indicating more extreme simplification, which is desired for this task. Similarly, the Llama-2-7b-chat-hf model demonstrated improvements with RL/RLHF, with the largest gains associated with the $R_{FKGL+OvS}$ reward (Exp. 20). In this configuration, RL (comparing between Exp. 18 and 20) increased SARI from 35.15 to 74.41, BLEU from 0.52 to 0.73, and ROUGE from 0.64 to 0.89. The $R_{FKGL+Rel}$ reward (Exp. 19) also resulted in improvements but to a lesser extent compared to the $R_{FKGL+OvS}$ reward (Exp. 20). In all experimentation, the FKGL score decreased when RL/RLHF was combined with SFT. The most prominent finding in the domain adaptation experimentation is that the best-performing domain-adapted LLM (Exp. 20) achieves similar performance to the best-performing in-domain models (Exp. 12). These findings underscore the potential of combining SFT with RL/RLHF to improve performance, presenting an approach for lowering annotation data requirements and cost.
@@ 4.3 Human Evaluation
The above assessment focuses on adequacy and fluency, where adequacy refers to the preservation of the original meaning, and fluency assesses the readability of the simplification. Automatic adequacy and fluency metrics provide useful information; however, these metrics do not holistically assess the quality of simplifications [75]. To provide a more comprehensive evaluation, we performed a manual evaluation of human- and AI-generated simplifications. The manual evaluation was carried out by three nurse practitioners with extensive knowledge of patient care and education, who were not involved in the SimpleDC annotation. The nurse practitioners reviewed a subset of the test samples in an A-B comparison, to evaluate the preservation of medical meaning and understandability for a broad audience. In this A-B comparison, the reviewers could indicate their preference for A, B, Both, or None. The reviewers were blinded to which samples, A or B, were generated by humans or AI, and the order (A vs. B assignments) was shuffled. The evaluation included: 1) human vs. the best performing Llama 2 (Exp. 20) and 2) human vs. the best performing GPT-4 (Exp. 17). The manual evaluation comprised a total of 104 sentence pairs, including 23 from CDC, 30 from NCI, and 51 from ACS.
Table 3 presents the results of the human evaluation, which demonstrate that human and Llama 2 simplifications consistently preserve the meaning of the original text and use understandable language. This further validates the results in Table 2, specifically that the trained Llama 2 model successfully learns the simplification task and emulates human simplification. The reviewers preferred human simplifications over GPT-4 simplifications to preserve meaning. The reviewers preferred the readability of GPT-4 over human simplification, suggesting a trade-off between fluency and adherence to the original content. This nuanced evaluation highlights the complexity of balancing adequacy and fluency in text simplification tasks, underscoring the need for a multifaceted approach in assessing model performance.
@@ 4.4 Error Analysis
Table 4: Error Analysis
Despite GPT-4’s advanced language capabilities and significantly larger model size, the trained Llama 2 models outperformed GPT-4 in both automated metrics and manual evaluations. This unexpected result can be attributed to differences in task interpretation. The simplification task was designed to improve the accessibility of the text for patients without altering its meaning. To better understand the editing performed by human annotators and LLMs to create the simplified versions, we calculated the proportion of tokens that are unchanged (no edits), inserted, and deleted from the original text. Table 4 presents a summary of the unchanged, inserted, and deleted words in the SimpleDC test set, normalized by the number of tokens in the original text. In creating SimpleDC, the annotators edited the NCI content more heavily than the ACS and CDC content because the NCI content started at a higher reading grade level (median 7.6 for ACS, 7.0 for CDC, and 8.3 for NCI). The Llama 2 model (Exp. 12) performed very light editing, deleting 9.3% of existing tokens and inserting 5.9% new tokens. In contrast, GPT -4 made many more edits, deleting 50.7% of existing tokens and inserting 61.4% new tokens. This discrepancy highlights the importance of task alignment between human expectations and AI model interpretations.
The GPT-4 performance was negatively impacted in the automatic and manual evaluations by over-generation. This tendency to add new text contrasts sharply with the more conservative approach of human annotators and the Llama 2 models, which introduced new content at rates of approximately 10.6% and 5.9% respectively, with average word counts of 14.24 and 14.05. Although these insertions can enhance understandability by adding explanatory content or simplifying complex ideas, they can also stray from the original message. For example, the GPT model simplified the sentence “The main functions of the liver include the following:” into a 30+ word sentence (“”The liver has some big jobs … help break down fat in food.”). While this type of generation may be helpful for the audience, it does not align with the objective of the task.
@@ 5 Discussion
Our research explores the effectiveness of LLMs for simplifying educational health information, contrasting the performance of SFT, RL, RLHF and in-context learning approaches. We uncover the essential role of model fine-tuning for domain-specific tasks, revealing how GPT models, despite their general prowess, require targeted adjustments to excel in specialized contexts. Central to our success is a novel reward model designed to optimize simplification by focusing on readability, relevance, and fidelity to the original text’s intent. This model incorporates metrics such as the Flesch-Kincaid Grade Level for readability and semantic similarity for content relevance, alongside an innovative Original vs. Simplified (OvS) reward informed by human feedback. Our findings demonstrate that our approach significantly improves the quality of text simplifications, making complex medical information more accessible and understandable to patients. The integration of human feedback into the reinforcement learning process, through the OvS reward, marks a significant advancement in aligning AI-generated content with human expectations. Overall, our study sets a new standard for employing AI in healthcare communication, highlighting the potential of finely tuned LLMs in bridging the gap between medical expertise and patient understanding.
The ability to accurately simplify complex medical texts without losing essential information is crucial for improving patient understanding and engagement. Our research illustrates the potential of using AI to assist in this process, but also the challenges in ensuring the simplifications are appropriate and accessible to all patients. This emphasizes the need for a collaborative approach involving technical innovation and input from healthcare professionals to ensure the simplifications meet the diverse needs of the patient population.
@@ 6 Conclusions
In this paper, we present SimpleDC, a carefully curated human-annotated text simplification corpus for digestive cancer education. The rich annotations of the dataset have been instrumental in training effective simplification models. We explore a range of learning approaches, including supervised fine-tuning, RL, RLHF, and prompt-based approaches. We introduced a new RLHF reward function, $R_{FKGL+Rel}$, which outperforms an existing RL reward function for text simplification. In terms of model performance, we observed that Llama 2 models, being supervised and data-dependent, outperformed GPT-4 in our experiments. This highlights the significant impact of high-quality domain-specific data in improving model performance. The addition of such data to Llama 2 models led to a substantial improvement in simplification quality, achieving a balance between comprehensibility and medical accuracy. Our results also demonstrate that RL/RLHF can be used to incorporate unlabeled text into training to improve simplification performance: 1) as a standalone training step, 2) in conjunction with supervised fine-tuning, and 3) as a domain adaptation technique. Although this represents a significant advancement, our findings also underscore the vast potential for future research in this area. Our work provides a robust foundation for further explorations, aiming to refine and expand the capabilities of NLP models to simplify complex medical information for enhanced patient understanding. Future research could explore integrating multimodal data sources, such as medical images or patient educational videos, alongside textual information to further enhance the simplification models’ ability to provide comprehensive and accessible health education. Additionally, investigating the incorporation of feedback mechanisms from actual patients and healthcare professionals into the model training process could significantly improve the relevance and usability of simplified medical content.
7 Research Data
The annotated corpus, SimpleDC, and all code will be made available upon acceptance of the manuscript.

### v3

@@ 4.1 Text Simplification
Table 3 presents the performance on the withheld SimpleDC test set, evaluating: Llama-2-7b-chat-hf, Llama-3-8B-Instruct, Llama-3-70B-Instruct, and GPT-4 Turbo. In the zero-shot setting (Exp. 01, 03, 05, and 15), Llama-2-7b-chat-hf achieved the highest SARI score, although GPT-4 achieved the lowest FKGL. SFT with the Llama models (Exp. 02, 04, and 06) markedly improved performance across SARI, BLEU, BERTScore, and ROUGE metrics, while reducing the FKGL by approximately two grade levels. These results underscore the role of SFT in understanding and learning the SimpleDC annotation schema and the annotators’ stylistic nuances. Notably, both Llama-2-7b-chat-hf and Llama-3-8B-Instruct showed substantial performance improvements with SFT, with 47.92 and 52.04 SARI and 7.08 and 6.32 FKGL, respectively, indicating effective task training. Llama-3-70B-Instruct also demonstrated noteworthy improvements with SFT (Exp. 06), achieving 46.73 SARI and the lowest reading level at 5.16 FKGL.
The success of the RL/RLHF varies by Llama variant and reward strategy. In RL/RLHF-only experiments, Llama-3-8B-Instruct (Exp. 11-13) consistently outperformed Llama-2-7b-chat-hf (Exp. 07-09) across multiple metrics, achieving higher SARI and lower FKGL. ForLlama-3-8B-Instruct, the proposed RLHF reward function, $R_{FKGL+OvS}$, markedly increased the SARI score from 20.44 (Exp. 11) in the RL scenario to 55.01 (Exp. 13). These findings emphasize the importance of integrating human feedback from the target task/data set through RLHF, as compared to the data set agnostic feedback of RL alone.
When combining SFT with RLHF for Llama-2-7b-chat-hf (Exp. 10), the model achieved the highest performance, at 60.39 SARI, 0.70 BLEU, and 0.97 BERTScore. This configuration also yielded a relatively low reading level at 8.24 FKGL, demonstrating that SFT and RLHF can be complementary in enhancing text simplification capabilities. However, for Llama-3-8B-Instruct, combining SFT with RLHF (Exp. 14) also showed no discernable improvement over RLHF alone (Exp. 13). In this configuration, the model achieved a SARI of 50.98, BLEU of 0.53, BERTScore of 0.91, and FKGL of 8.11.
For in-context learning, GPT-4 Turbo (Exp. 15-19) achieved lower scores across all adequacy metrics (SARI, BLEU, BERTScore, and ROUGE) compared to the Llama models (Exp. 01-14) but achieved the lowest FKGL. Lower performance does not necessarily indicate lower quality simplifications generated by GPT-4. Based on our qualitative review, GPT-4 edited the original text more extensively, resulting in a larger divergence between the original and simplified text. Additionally, the voice or writing style of GPT-4 appears to differ from that of our annotators, contributing to the lower performance for adequacy metrics. The FKGL-Enhanced prompting strategy (Exp. 17) achieved the best performance across adequacy metrics, and the CoT prompting strategy (Exp. 18) achieved the lowest FKGL. The Self-Correction approach (Exp. 19) achieved a balance between adequacy and reading level at 22.86 SARI and 4.48 FKGL.
@@ 4.2 Domain Adaptation
We used the Med-EASi corpus to explore the role of RL/RLHF in domain adaptation by first conducting out-of-domain SFT with Med-EASi, followed by in-domain RL/RLHF training using unlabeled digestive cancer text. Table 3 (Exp. 20-25) presents the domain adaptation results. This experimentation evaluated the feasibility of leveraging out-of-domain corpora to enhance in-domain performance. In this domain adaptation setting, the performance gains from RL/RLHF vary by Llama variant and specific reward function. For Llama-3-8B-Instruct, continued in-domain RLHF using the $R_{FKGL+OvS}$ reward markedly improved performance over out-of-domain SFT on Med-EASi alone, increasing SARI from 38.73 to 51.06, BLEU from 0.43 to 0.57, and ROUGE from 0.60 to 0.72 (Exp. 23 vs. Exp. 25). However, Llama-3-8B-Instruct with RLHF alone and no prior SFT on Med-EASI (Exp. 13) achieved better results overall, indicating the combined out-of-domain SFT and in-domain RLHF training were not complementary.
For Llama-2-7b-chat-hf, RL/RLHF alone (Exp. 07-09) was not as effective; however, the best performance across all models was achieved by combing out-of-domain SFT on Med-EASi with in-domain RLHF with the $R_{FKGL+OvS}$ reward. For this configuration, continued training through in-domain RLHF increased SARI from 35.15 to 74.41, BLEU from 0.52 to 0.73, and ROUGE from 0.64 to 0.89 and decreased FKGL from 8.49 to 7.74 (Exp. 20 vs Exp. 22). The $R_{FKGL+Rel}$ reward (Exp. 21) also led to improvements over out-of-domain SFT alone but was less effective compared to the $R_{FKGL+OvS}$ reward. The domain adaptation results highlight that the top-performing domain-adapted LLM (Exp. 22) achieves comparable performance to the best in-domain models (Exp. 10, 13, 04 and 14). These findings affirm the potential of combining SFT with RL/RLHF to enhance performance while minimizing annotation data requirements and associated costs.
@@ 4.3 Human Evaluation
The above assessment focuses on adequacy and fluency, where adequacy refers to the preservation of the original meaning, and fluency assesses the readability of the simplification. Automatic adequacy and fluency metrics provide useful information; however, these metrics do not holistically assess the quality of simplifications [86]. To provide a comprehensive evaluation, we manually evaluated human- and AI-generated simplifications. Manual evaluations were performed by three nurse practitioners with extensive knowledge of patient care and education, who were not involved in the SimpleDC annotation. The nurse practitioners reviewed a subset of the test samples in an A-B comparison, to evaluate the preservation of medical meaning and understandability for a broad audience. In this A-B comparison, the reviewers could indicate their preference for A, B, Both, or None. The reviewers were blinded to which samples, A or B, were human- or AI-generated, and the order (A vs. B assignments) was shuffled. The evaluation included: 1) human vs. the best performing Llama (Exp. 22) and 2) human vs. the best performing GPT-4 (Exp. 19). The manual evaluation comprised a total of 104 sentence pairs, including 23 from CDC, 30 from NCI, and 51 from ACS.
Table 4 presents the results of the human evaluation, which demonstrate human and Llama simplifications consistently preserve the meaning of the original text and use understandable language. This further validates the results in Table 3, specifically that the trained Llama model successfully learns the simplification task and emulates human simplification. The reviewers preferred human simplifications over GPT-4 simplifications to preserve meaning. The reviewers preferred the readability of GPT-4 over human simplification, suggesting a trade-off between fluency and adherence to the original content. This nuanced evaluation highlights the complexity of balancing adequacy and fluency in text simplification tasks, underscoring the need for a multifaceted approach in assessing model performance.
@@ 4.4 Error Analysis
Table 5: Error Analysis of the best Llama 2 model (Exp 22) and the best GPT-4 model (Exp 19)
Despite GPT-4’s advanced language capabilities and significantly larger model size, the trained Llama models outperformed GPT-4 in both automated metrics and manual evaluations. This unexpected result can be attributed to differences in task interpretation. The simplification task was designed to improve the accessibility of the text for patients without altering its meaning. To better understand the editing performed by human annotators and LLMs to create the simplified versions, we calculated the proportion of tokens that are unchanged (not edited), inserted, and deleted from the original text. Table 5 presents a summary of the unchanged, inserted, and deleted words in the SimpleDC test set, normalized by the number of tokens in the original text (comparing with the Llama 2 model outputs from Exp. 22). In creating SimpleDC, the annotators edited the NCI content more heavily than the ACS and CDC content because the NCI content started at a higher reading grade level (median 7.6 for ACS, 7.0 for CDC, and 8.3 for NCI). The Llama 2 model performed very light editing, deleting 9.3% of existing tokens and inserting 5.9% new tokens. In contrast, GPT-4 made many more edits, deleting 50.7% of existing tokens and inserting 61.4% new tokens. This discrepancy highlights the importance of task alignment between human expectations and AI model interpretations.
The GPT-4 performance was negatively impacted in the automatic and manual evaluations by over-generation. This tendency to add new text contrasts sharply with the more conservative approach of human annotators and the Llama models, which introduced new content at rates of approximately 10.6% and 5.9% respectively, with average word counts of 14.24 and 14.05. Although these insertions can enhance understandability by adding explanatory content or simplifying complex ideas, they can also stray from the original message. For example, the GPT model simplified the sentence “The main functions of the liver include the following:” into a 30+ word sentence (“The liver has some big jobs … help break down fat in food.”). While this type of generation may be helpful for the audience, it does not align with the objective of the task.
@@ 5 Discussion
Our research explores the effectiveness of LLMs for simplifying educational health information, contrasting the performance of SFT, RL, RLHF, and in-context learning approaches. We explore the role of model training for specific tasks, revealing how GPT models, despite their general prowess, require targeted adjustments to excel in specialized contexts. We introduce a novel reward model designed to optimize simplification by focusing on readability, relevance, and fidelity to the original text’s intent. This new model incorporates an automatic reading-level metric (FKGL) and the innovative Original vs. Simplified (OvS) reward informed by human feedback. Our findings demonstrate how RLHF can improve the quality of text simplifications, making complex medical information more accessible and understandable to patients. The integration of human feedback into the reinforcement learning process, through the OvS reward, marks a significant advancement in aligning AI-generated content with human expectations. Our study advances AI-driven text simplification for healthcare communication and demonstrates the potential for task-specific LLMs to bridge the gap between medical expertise and patient understanding.
Our results indicate that foundational LLMs, with their unique architectures and training protocols, respond differently to continued training. Specifically, Llama-2 and Llama-3 exhibited varied responses to SFT and RLHF. For Llama-2, SFT improved performance across all metrics compared to zero-shot; however, RLHF did not significantly alter performance relative to zero-shot. Combining SFT and RLHF for Llama-2 resulted in improved performance relative to either method used individually in both in-domain and domain adaptation experiments. In contrast, for Llama-3, both SFT and RLHF independently improved performance, with in-domain RLHF training yielding the best results. However, combining SFT and RLHF for Llama-3 did not enhance performance beyond that of RLHF alone. The architectural and training specifics for Llama-3 are not yet available, making it difficult to fully understand the differences between Llama-2 and Llama-3 and the reasons for these performance discrepancies.
The ability to accurately simplify complex medical texts without losing essential information is crucial for improving patient understanding and engagement. Our research illustrates the potential of using AI to assist in this process but also highlights the challenges in ensuring the simplifications are appropriate and accessible to all patients. This emphasizes the need for a collaborative approach that combines technical innovation with input from healthcare professionals to ensure the simplifications meet the diverse needs of the patient population. By leveraging AI’s capabilities alongside expert human feedback, we can better tailor health information to enhance patient comprehension and empowerment.
@@ 6 Conclusions
In this paper, we present SimpleDC, a carefully curated human-annotated text simplification corpus for digestive cancer education. The rich annotations of the dataset have been instrumental in training effective simplification models. We explore a range of learning approaches, including SFT, RL, RLHF, and prompt-based approaches. We introduced a new RLHF reward function, $R_{FKGL+OvS}$, which outperforms an existing RL reward function for text simplification. In terms of model performance, we observed that Llama models with task-specific training outperformed GPT-4. This highlights the impact of high-quality task-specific annotated data in improving model performance and simplification quality, including balancing comprehensibility and medical accuracy. Our results also demonstrate that RL/RLHF can be used to incorporate unlabeled text into training to improve simplification performance: 1) as a standalone training step, 2) in conjunction with SFT, and 3) as a domain adaptation technique. Although this represents a significant advancement, our findings also underscore the vast potential for future research in this area. Our work provides a robust foundation for further explorations, aiming to refine and expand the capabilities of NLP models to simplify complex medical information for enhanced patient understanding. Future research could explore integrating multimodal data sources, such as medical images or patient educational videos, alongside textual information to further enhance the simplification models’ ability to provide comprehensive and accessible health education. Additionally, investigating the incorporation of feedback mechanisms from actual patients and healthcare professionals into the model training process could significantly improve the relevance and usability of simplified medical content.
7 Data Availability and Legal Considerations
We provide the new SimpleDC corpus to the research community to advance health text simplification. [^6] https://github.com/mushfiqur11/simpledc-dataset.git In developing SimpleDC, we utilized publicly available text data from the NCI, ACS, and CDC websites with the goal of advancing text simplification and public health communication. Our use of this data is in alignment with fair use principles, as the data was employed exclusively for academic and research purposes, without any commercial intent. We minimized data usage, extracting only the portions necessary to support our work on text simplification and public health communication. Furthermore, our use of the data does not adversely affect the market value or potential of the original content. The text data from the NCI, ACS, and CDC was specifically used to improve NLP methods that enhance the accessibility of health information, with all sources carefully attributed. SimpleDC is made publicly available, with detailed source attribution, ensuring full transparency and traceability.
@@ 7 Data Availability and Legal Considerations
Appendix A Case Study of Simplification Outputs
Source/Model
Exp.
No.
Simplification
Output
Original - As food is digested, it passes from the stomach to the small intestine.
Target Simplification - As food is broken down in your body, it passes from the stomach to the section of the gut called the small intestine.
Llama-2-7b with SFT(SimpleDC) + RLHF (SimpleDC) 10 Food is digested as it passes from the stomach to the small intestine.
Llama-2-7b with SFT(MedEASi) + RLHF (SimpleDC) 22 At this point, food passes from the stomach to the colon.
Llama-3-8b with SFT(SimpleDC) + RLHF (SimpleDC) 14 Food is digested and it passes from the stomach to the small intestine.
Llama-3-8b with SFT(MedEASi) + RLHF (SimpleDC) 25 As food is digested, it goes from the stomach to the small intestine.
Llama-3-70b with SFT(SimpleDC) 06 When you eat food, it goes from the stomach to the small intestine. This is called the small bowel. It is like a long, thin, winding road.
GPT-4 with Self-correction 19 When we eat, the food goes from our stomach to our small intestine.
Original - Basal cell carcinomas are a type of skin cancer that can develop in the perianal skin.
Target Simplification - Basal cell cancer is a kind of skin cancer that can sometimes start in the perianal skin.
Llama-2-7b with SFT(SimpleDC) + RLHF (SimpleDC) 10 Basal cell carcinomas are a type of skin cancer that can develop in the perianal skin.
Llama-2-7b with SFT(MedEASi) + RLHF (SimpleDC) 22 Basal cell cancer is a type of skin cancer that can develop in the perianal skin.
Llama-3-8b with SFT(SimpleDC) + RLHF (SimpleDC) 14 Basal cell cancers are a type of skin cancer that can develop in the perianal skin.
Llama-3-8b with SFT(MedEASi) + RLHF (SimpleDC) 25 Basal cell carcinomas are a type of skin cancer that can develop in the perianal skin.
Llama-3-70b with SFT(SimpleDC) 06 Basal cell carcinomas are a kind of skin cancer that can grow in the perianal skin. This is called anal melanosis.
GPT-4 with Self-correction 19 Basal cell carcinomas are a kind of skin cancer. They can grow on the skin around the bottom.
Table 6: Simplification outputs of the same sentence by different models.
Table 6 presents text simplification examples from Llama-2 and Llama-3 models using various fine-tuning and reinforcement learning approaches. Both Llama-2-7b and Llama-3-8b models generally produce concise and accurate simplifications. However, the Llama-3-70b model, while generating seemingly more readable outputs, tends to add additional information not present in the original text. These results underscore the importance of model choice and configuration in text simplification, highlighting the need for precise model selection and fine-tuning in domain-specific applications.